Observer: Arcidiacono,C, Boutsia, K, (INAF), D.Miller (LBTO)
Telescope Operator:Bechetti G
SW Support:
AO Support: D.Miller
Telescope Support: Kevin
Instrument Support: Edwards,M


06:48UT we close for winds and we stayed closed for the rest of the night...

Project_RUN Objects Filter ExpTime Notes
Cusano 652345526 Ks/J 530s,540s COMPLETED - the loop was quite unstable during this source - we had to repeat preset several times
Cusano 659717590 Ks/J 300s,1200s COMPLETED - the second offset for J was for a really low elev
Calib M15 all filters   this field was obtained for plate scale calibration and for the orientation of the detector - necessary calib. for all astrometry projects
Antoniucci PVCep ~ ~ we tried a preset but the target is very faint in the R band so the AO loop was unstable and we did not manage to proceed


00:45 we tried to obtain Ks flats but it was too late - the same for J.
  • we take some short darks: 15s,10s,5s,3s,2s

01:24 the TO is collimating

01:47 slew to Cusano 652345526
  • the preset failed twice - there might be something wrong with the pointing

01:52 TO corrects pointing
  • we preset to the target again
  • the star was collimated - but once the AO took over, the star lost collimation
  • not off-loading tip/tilt ?? - restart AOS??
  • the AO failed eventually

02:05 preset to Cusano again using a different guide star.
  • the probe was vignetting the previous one
  • preset successful this time!

02:15 the AO loop opened during the offset
  • send preset again

02:21 dither pattern in KJJK
  • we checked the focus for the K filter and it is the correct one.
  • magnitude on the sensor R=13.5
  • it does not off-load the high modes!! - WHY???

02:28 The loop is open!!
  • we try restarting the AOS!!!
  • there is a huge amount of Z11 on the active optics
  • doug is suggesting to clear active optics in both primary and secondary

02:33 send the preset again
  • the active optics gave numbers that indicate collimation but in was not TRUE
  • the AO loop failed
  • we asked the TO to collimate on a nearby star
  • the GCS is having trouble picking up the correct star to collimate

02:53 slew back to the target
  • turn mirror ventilation off

02:58 send dither pattern again

03:09 the AO loop is open

03:15 send preset again
  • start dither pattern again JK,

03:31 Adjust the CCD47 level in order to better see the AO reference star

03:33 slew to Cusano - 659717590
  • preset went through
  • magnitude on the guider R=13.8

03:41 dither pattern KJJK
  • we only did KJJ - we aborthed the last Ks dither pattern because elev~30deg

04:15 slewing to M15 for the plate scale
  • Alfio is online helping with the 'click'n'go' procedure
  • the guide star has a lot of astigmatism - go to collimate on a nearby star

04:24 we are collimated - send the preset again

04:35 we start M15 dither pattern in all filters
  • North is on the right and east is down on the Pisces image for PA=0
  • DIMM - 1.2"

05:17 offsetting for the sky

05:46 finished with the PlateScale/Orientation calibration 15min earlier
  • we slew to PVCep to get some images and judge feasibility
  • mag on the AO detector R=15.6 - remember to find PSF with R=15.5
  • AO skip frames - the AO might fail ... we have no wind info !!! - it was reset by DLM and it works
  • it has problems to close the AO loop
  • we have 0.6 mag di assorbimento - cirrus on the allsky camera and the humidity is climbing
  • the target is really faint in the R band so even slight seeing variations or clouds result on a variable AO correction
  • difficult to obtain results on this source - in addition clouds are accumulating

06:08 slew to Cusano V1176_Tau which is brighter
  • we need to unwrap the telescope

06:20 Thick clouds! - we cannot find guide stars

06:30 slew to Cusano V1176_Tau - we still have clouds!!
  • error from the AO side - at the beginning the star was reported at R=11.0 but while it was correcting it went down to R=18.
  • the star is variable due to clouds passing through
  • we can not find the star - waiting for the clouds to clear out

06:48 CLOSE DOME for winds
  • the winds are getting close to critical - clouds keep building - the forecast is not very good
  • the TO decides to close the dome for now and he will keep an eye in case conditions get better

Cut and Paste of the IDL command line

Observe commands and IDL console in the following:


-- KonstantinaBoutsia - 13 Oct 2013
Topic revision: r6 - 15 Oct 2013, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback